OpenStudy (anonymous):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book?

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

20.7

OpenStudy (dangerousjesse):

I meant \(18\times .15=x\), \(18-x=y\), \(y\times 1.06= the~ answer\)

OpenStudy (anonymous):

1.06y

OpenStudy (anonymous):

i think

OpenStudy (dangerousjesse):

What's \(y\)?

OpenStudy (anonymous):

im confused on this problem

OpenStudy (dangerousjesse):

First, just solve for x. \(18\times.15=x\)

OpenStudy (dangerousjesse):

That's how you'll find out how much is discounted.

OpenStudy (anonymous):

thats 2.7

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Now multiply 15.3 (\(-2.7+18\)) by 1.06 to find your answer :)

OpenStudy (anonymous):

16.218

OpenStudy (dangerousjesse):

Because there's no such thing as $16.218, round that up to the nearest hundredth to find your answer :)